need help with e,f,g Let W be the weight of the ith airline passengers luggage. Assume that the weights are independent, cach with pdf f(w) - B-w
Answer #1

Given that +(W) = OB° wol it 0 LWLB 30, +(W) = 2 WO; OLWLB €(W) = o swodw +1 var(w) - Scotdw-eoo = 0². 0 loto coto Bio Fotb][100 T 80² x 3/5x12 - P(Y>0 1613) -1- (0.1613) Diriwn -0/28) P(unse) = P(w. ---Wn>) ={P(WIDE). Lo worldw] [reso -{{ e. 10-{-(0so, By defination, wnn. Ly B d het u = (wan) p(4=y) - pl won symm) =P(wona Byln) (Pl. 2 By n)]? - BY 417)®_y® clearly Oxy</